# Visual Odometry with Transformers

### v1

@@ Abstract
Modern monocular visual odometry methods typically combine pre-trained deep learning components with optimization modules, resulting in complex pipelines that rely heavily on camera calibration and hyperparameter tuning, and often struggle in unseen real-world scenarios. Recent large-scale 3D models trained on massive amounts of multi-modal data have partially alleviated these challenges, providing generalizable dense reconstruction and camera pose estimation. Still, they remain limited in handling long videos and providing accurate per-frame estimates, which are required for visual odometry. In this work, we demonstrate that monocular visual odometry can be addressed effectively in an end-to-end manner, thereby eliminating the need for handcrafted components such as bundle adjustment, feature matching, camera calibration, or dense 3D reconstruction. We introduce VoT, short for Visual odometry Transformer, which processes sequences of monocular frames by extracting features and modeling global relationships through temporal and spatial attention. Unlike prior methods, VoT directly predicts camera motion without estimating dense geometry and relies solely on camera poses for supervision. The framework is modular and flexible, allowing seamless integration of various pre-trained encoders as feature extractors. Experimental results demonstrate that VoT scales effectively with larger datasets, benefits substantially from stronger pre-trained backbones, generalizes across diverse camera motions and calibration settings, and outperforms traditional methods while running more than $3\times$ faster. The code will be released.
@@ 1 Introduction
The goal of monocular visual odometry is to estimate a camera’s position and orientation from a sequence of video frames (Cadena et al., 2016). In recent years, monocular visual odometry has gained increasing attention across various fields, including augmented reality, virtual reality, autonomous driving, and robotics. This growing interest is particularly notable when compared to systems that rely on stereo vision (Wang et al., 2017a; Engel et al., 2014) or multimodal inputs, such as visual-inertial odometry (Von Stumberg et al., 2018; Forster et al., 2015). Although monocular setups present greater challenges, their simplicity in deployment enables even broader applicability in real-world scenarios. Therefore, this paper focuses on monocular visual odometry using neural network-based approaches.
@@
There has been rapid progress in monocular visual odometry over a short period, highlighting the potential of learning-based approaches. The dominant approach brought by this progress utilizes learnable features as well as differentiable optimization layers within deep learning frameworks (Chen et al., 2024; Teed et al., 2023). These designs improve convergence and robustness in real-world conditions. However, their performance critically relies on postprocessing techniques such as bundle adjustment or feature matching to refine the estimates of camera poses. The use of such traditional ‘feature-engineered’ components introduces hand-crafted elements and task-specific hyperparameters to incorporate prior domain knowledge. While proven effective, these practices limit the scaling behavior of models as they do not fully exploit the end-to-end nature of deep learning systems (Tay et al., 2023). Additionally, auxiliary supervision (e.g., by dense optical flow) is often used to support training, further increasing system complexity and limiting training data. Motivated by these limitations, we propose a direct ‘end-to-end’ pose regression approach that eliminates hand-crafted modules and auxiliary tasks.
@@
Our method draws inspiration from recent successes in end-to-end structured prediction tasks, such as machine translation (Vaswani et al., 2017), image recognition (Dosovitskiy et al., 2021), and object detection (Carion et al., 2020). By learning a mapping directly from video input to camera poses, the model becomes less dependent on predefined priors and can better capture complex temporal and spatial relationships. This design choice aligns with recent evidence (Nguyen et al., 2025) suggesting that reducing inductive bias enables models to generalize more effectively from data. Indeed, recent breakthroughs in image recognition (Dosovitskiy et al., 2021), 2D/3D object detection (Carion et al., 2020; Nguyen et al., 2022), and 3D reconstruction (Wang et al., 2025a; Leroy et al., 2024) increasingly adopt Transformer-based architectures for learning representations with large-scale datasets. We hypothesize that an end-to-end visual odometry framework with a transformer is a key enabler for scalability, leading to improved robustness and accuracy.
In this paper, we introduce the Visual odometry Transformer (VoT), a fully end-to-end framework for monocular visual odometry. Our approach formulates visual odometry as a direct pose prediction task, using a transformer-based encoder-decoder architecture. The encoder is initialized with a pre-trained backbone, while the decoder employs both temporal and spatial attention to model interactions across frames. This design implicitly captures motion-specific properties, such as feature matching, without requiring explicit supervision or auxiliary modules.
In our framework, the VoT predicts the sequence of relative camera poses in a single forward pass and is trained end-to-end using loss functions for both translation and rotation. To ensure the validity of rotation predictions, we project them onto the SO(3) manifold and employ a loss function that computes the shortest distance between the predicted and ground-truth rotation matrices. Importantly, our implementation avoids any custom CUDA operations or non-standard layers, improving reproducibility and compatibility with common deep learning frameworks.
We train and evaluate VoT on several large-scale indoor and outdoor datasets where our method achieves competitive accuracy across all metrics. In addition, VoT shows strong scaling behavior and can generalize to unseen datasets. We find that spatial-temporal attention enables the model to implicitly learn global feature interactions, supporting our hypothesis that an end-to-end design trained on extensive data can outperform systems that rely heavily on hand-crafted priors and post-processing. Notably, VoT with the end-to-end design runs more than $3\times$ faster than its counterparts.
@@ 2 Related Work
Visual odometry. Visual odometry systems estimate the position and orientation of a camera using video input. Unlike SLAM which corrects errors via loop closure (Cadena et al., 2016; Campos et al., 2021; Yugay et al., 2024), these systems tend to accumulate camera tracking errors (drift). Many different modalities of visual odometry have been explored by past work, including visual-inertial odometry (Forster et al., 2015; Von Stumberg et al., 2018) and stereo visual odometry (Engel et al., 2014; Wang et al., 2017a). Here, we focus on the monocular case, where the only input is a monocular video stream. Traditional monocular visual odometry methods (Engel et al., 2014; 2018; Campos et al., 2021) are sensitive to illumination and rolling shutter artifacts, and more importantly, cannot adequately estimate the scale of the scene. To overcome these limitations, VoT leverages a large pre-trained encoder and extensive training data for robust generalization.
@@
Deep monocular visual odometry. Deep learning has advanced monocular visual odometry in both supervised (Wang et al., 2017b; 2020; Teed & Deng, 2020a; Teed et al., 2023) and unsupervised (Yin & Shi, 2018; Ranjan et al., 2019; Sharma & Ventura, 2019; Li et al., 2020) settings. DeepVO (Wang et al., 2017b) uses recurrent networks for temporal modeling, while SfMLearner (Sharma & Ventura, 2019) jointly learns depth and motion without labels. DPVO (Teed et al., 2023), inspired by RAFT (Teed & Deng, 2020b), combines flow, confidence, and geometry with iterative GRU updates, while LeapVO (Chen et al., 2024) increases robustness in dynamic scenes via keypoints.
@@
Prior work introduced TSFormer (Françani & Maximo, 2025), which predicts camera poses directly from video sequences using an end-to-end approach. However, its performance remains below that of hand-crafted methods (Teed et al., 2023; Chen et al., 2024) due to architecture, small training dataset, and rotation representation. Existing methods generalize poorly because they rely on small, hard-to-scale architectures, complex hand-crafted components, and often require camera parameters, limiting real-world use. By comparison, VoT uses a scalable end-to-end design that works without camera parameters, enabling deployment across diverse scenarios.
Large 3D models. A trend in 3D vision is the development of large-scale models that jointly estimate camera poses and dense geometry (Wang & Agapito, 2025; Wang et al., 2025a; b), or learn versatile representations for downstream tasks (Wang et al., 2024; Leroy et al., 2024). These models are increasingly viewed as foundational, yet they suffer from drift on long video sequences and produce scale-ambiguous poses requiring calibration. In contrast, VoT focuses solely on camera pose estimation, showing improved accuracy and less drift. Moreover, our model is significantly faster.
@@
Video transformers. Since their introduction in machine translation (Vaswani et al., 2017), transformers have become the dominant architecture in NLP (Devlin et al., 2019; OpenAI, 2023) and vision (He et al., 2022; Oquab et al., 2024). Central to this framework, self-attention aggregates information across entire sequences, but its quadratic complexity makes direct application to videos inefficient. To mitigate this, prior work has proposed factorizing 3D attention into spatial and temporal components (Zhang et al., 2021) or employing efficient variants such as divided space-time, sparse local-global, and axial attention (Bertasius et al., 2021; Arnab et al., 2021; Zhao et al., 2022). In this work, we adopt the principle of efficient video learning by decoupling self-attention into separate temporal and spatial attention components.
@@ 3 Method
In this section, we describe the architecture of VoT; see Fig.˜1 for its overview. It consists of three main components: a pre-trained encoder to extract a compact representation of each frame, a transformer-based decoder, and a small feed-forward network that makes the final pose predictions.
Unlike previous methods such as DVPO (Teed et al., 2023) or LeapVO (Chen et al., 2024), VoT can be implemented in any deep learning framework, thanks to its streamlined architecture. Following the spirit of end-to-end frameworks (Carion et al., 2020), our inference code can be written purely in PyTorch, without additional complexities. We believe that the simplicity and accessibility of our approach will encourage further participation from the research community.
@@ 3.1 Architecture
Pre-trained encoder. We utilize pre-trained networks commonly employed in previous studies (Wang et al., 2024; 2025b) for encoding and extracting features from video frames. As the Vision Transformer (ViT) (Dosovitskiy et al., 2021) has become the dominant architecture in computer vision, our pre-trained encoders are based on ViT. Specifically, we denote the sequence of input video frames as $V\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ is the number of frames in the video and $H,W$ denote the resolution of each frame. Each input image is then tokenized into $(h\cdot w)$ non-overlapping patches with $h{=}\frac{H}{p},w{=}\frac{W}{p}$ and $p$ as patch size. These tokens are processed through a series of ViT layers, producing image features $F\in\mathbb{R}^{T\times(h\cdot w)\times d}$, where $d$ is the hidden dimension of the ViT. The ViT layer follows the standard architecture, consisting of a multi-head self-attention module and a feed-forward network. Since the transformer architecture is permutation-invariant, we augment it with sinusoidal position encodings (Vaswani et al., 2017).
@@
$\hat{Q}=\hat{K}=\hat{V}=F\text{[:, 1:, :]}^{\top};\qquad\qquad\,\in\mathbb{R}^{(h\cdot w)\times T\times d}\\$ (1)
$\hat{\mathrm{head}}_{i}=\mathrm{Attention}(\hat{Q}W^{\hat{Q}}_{i},\hat{K}W^{\hat{K}}_{i},\hat{V}W^{\hat{V}}_{i}),\quad\in\mathbb{R}^{(h\cdot w)\times T\times d_{h}}\\$ (2)
@@
$\hat{F}=\mathrm{Concat}(\hat{\mathrm{head}}_{i},...,\hat{\mathrm{head}}_{h})^{\top}W^{\hat{O}},\qquad\qquad\,\,\in\mathbb{R}^{T\times(h\cdot w)\times d}\\$ (3)
$\mathrm{TemporalAttention}(\hat{Q},\hat{K},\hat{V})=[F\text{[:,:1,:]},\hat{F}],\qquad\qquad\,\,\in\mathbb{R}^{T\times(h\cdot w+1)\times d}\\$ (4)
@@
$\bar{Q}=\bar{K}=\bar{V}=\hat{F}\qquad\qquad\,\in\mathbb{R}^{T\times(h\cdot w+1)\times d_{h}}\\$ (5)
$\bar{\mathrm{head}}_{i}=\mathrm{Attention}(\bar{Q}W^{\bar{Q}}_{i},\bar{K}W^{\bar{K}}_{i},\bar{V}W^{\bar{V}}_{i})\quad\in\mathbb{R}^{T\times(h\cdot w+1)\times d_{h}}\\$ (6)
@@
$\mathrm{SpatialAttention}(\bar{Q},\bar{K},\bar{V})=\mathrm{Concat}(\bar{\mathrm{head}}_{i},...,\bar{\mathrm{head}}_{h})W^{\bar{O}}\qquad\,\in\mathbb{R}^{T\times(h\cdot w+1)\times d}\\$ (7)
@@
At the end, the camera embedding, $e\in\mathbb{R}^{T\times d}$, will be used to predict the camera poses for each frame.
@@ 3.2 Camera Poses Prediction
To regress the relative camera poses, the regression head takes a camera embedding from each frame as its input. A single linear layer then projects the embeddings to the rotation $\mathbb{F}_{R}\in\mathbb{R}^{(T-1)\times 9}$ and translation $\mathbb{F}_{t}\in\mathbb{R}^{(T-1)\times 3}$ vectors. Rotation vectors are projected to the closest valid rotation matrix by solving the special orthogonal Procrustes (Brégier, 2021) problem:
@@
The optimization problem is solved via singular value decomposition of $\mathbb{F}_{R}$ following Umeyama (1991). $\mathbb{F}_{t}\in\mathbb{R}^{(T-1)\times 3}$ is used as a translation prediction directly.

### v2

@@ Abstract
Despite the rapid development of large 3D models, classical optimization-based approaches dominate the field of visual odometry (VO). Thus, current approaches to VO heavily rely on camera parameters and many hand-crafted components, most of which involve complex bundle adjustment and feature matching processes. Although disregarded in the literature, we find it problematic in terms of both (1) speed, that performs bundle adjustment requires a significant amount of time; and (2) scalability, as hand-crafted components struggle to learn from large-scale training data. In this work, we introduce a simple yet efficient architecture, Visual odometry Transformer (VoT), that formulates monocular visual odometry as a direct relative pose regression problem. Our approach streamlines the monocular visual odometry pipeline in an end-to-end manner, effectively eliminating the need for handcrafted components such as bundle adjustment, feature matching, or camera calibration. We show that VoT is up to $\sim 4$ times faster than traditional approaches, yet with competitive or better performance. Compared to recent 3D foundation models, VoT runs $\sim 10$ times faster with strong scaling behavior in terms of both model sizes and training data. Moreover, VoT generalizes well in both low-data regimes and previously unseen scenarios, reducing the gap between optimization-based and end-to-end approaches. Code will be released.
@@ 1 Introduction
The goal of monocular visual odometry is to estimate a camera’s position and orientation from a sequence of video frames [cadena2016past_present_future_of_slam]. In recent years, monocular visual odometry has gained increasing attention across various fields, including augmented reality, virtual reality, autonomous driving, and robotics. This growing interest is particularly notable when compared to systems that rely on stereo vision [wang2017stereo_dso, engel2014lsd] or multimodal inputs, such as visual-inertial odometry [stumberg2018direct_sparse_visual_inertial, christian2015robotics_science_systems]. Although monocular setups present greater challenges, their simplicity in deployment enables even broader applicability in real-world scenarios. Therefore, this paper focuses on monocular visual odometry using neural network-based approaches.
@@
There has been rapid progress in monocular visual odometry over a short period, highlighting the potential of learning-based approaches. The dominant approach brought by this progress utilizes learnable features as well as differentiable optimization layers within deep learning frameworks [chen2024leapvo, teed2023dpvo]. While these designs improve convergence and robustness in real-world conditions, their performance critically relies on post-processing techniques, such as bundle adjustment or feature matching, to refine the estimates of camera poses. However, the limitations are still evident in real-world applications as the prediction in the wild has to undergo a slow post-processing process. Furthermore, hand-crafted elements incorporate prior domain knowledge, making it hard to scale with training data.
@@
To this end, we shift our attention to the fast and scalable architecture. Recent successes in machine translation [vaswani2017transformer], image recognition [dosovitskiy2020vit], and object detection [carion2020detr] demonstrated that a simple transformer is effective enough for end-to-end structured prediction. By learning a mapping directly from video input to camera poses, the model becomes less dependent on predefined priors and can better capture complex temporal and spatial relationships. This design choice aligns with recent evidence [nguyen2025pit], suggesting that reducing inductive bias enables models to generalize more effectively from data. Most importantly, the end-to-end formulation enables direct prediction of camera trajectories without relying on heavy post-processing.
In this paper, we introduce the Visual odometry Transformer (VoT), a fully end-to-end framework for monocular visual odometry. Our approach formulates visual odometry as a direct pose prediction task. It mainly differs from previous models in the use of encoder-decoder architecture. The encoder is initialized with a pre-trained backbone to extract image features, while the decoder employs both temporal and spatial attention to model interactions across frames. The decouple between feature extraction with encoder and feature interaction with decoder significantly boosts the inference speed and scalability of the model. Fig. 1 shows that our end-to-end model is around ten times faster than 3D foundation models and at least three times than those with hand-crafted components while exhibiting similar or even better performance on multiple benchmarks.
Our key contributions can be summarized as follows:
  • VoT is the simple architecture by far for visual odometry as it commissions the pre-trained transformer as encoder and a light-weight transformer as decoder. This design inherently leads to significant speedup in our runtime while still being scalable in terms of both model sizes and training data;
  • For the first time, we achieve competent performance on several benchmarks of visual odometry without using any hand-crafted components and simply predicting camera poses in an end-to-end manner;
  • We showcase that end-to-end learning with VoT is beneficial especially when large-scale train data is available while it still generalizes well to low-regime data and unseen scenarios.
@@ 2 Related Work
Visual odometry. Visual odometry systems estimate the position and orientation of a camera using video input. Unlike SLAM, which corrects errors via loop closure [cadena2016past_present_future_of_slam, campos2021orbslam3, yugay2024magicslam], these systems have to cope with camera tracking errors (drift). Many different modalities of visual odometry have been explored by past work, including visual-inertial odometry [christian2015robotics_science_systems, stumberg2018direct_sparse_visual_inertial] and stereo visual odometry [engel2014lsd, wang2017stereo_dso]. While improving the overall performance, the multi-modal setting limits the potential to scale up the system with training data. Here, we focus on the monocular case, where the only input is a monocular video stream. Traditional monocular visual odometry methods [engel2014lsd, engel2018direct, campos2021orbslam3] are sensitive to illumination and rolling shutter artifacts, and more importantly, cannot adequately estimate the scale of the scene. To overcome these limitations, VoT leverages a large pre-trained encoder and extensive training data for robust generalization.
@@
Deep monocular visual odometry. Deep learning has further advanced monocular visual odometry in both supervised [wang2017deepvo, tartanvo2020tartanVO, teed2020deepv2d, teed2023dpvo] and unsupervised [yin2018geonet, ranjan2019compeptitivecollaboration, alisha2019unsupervised_learning_of_depth_and_ego_motion, li2020selfsuperviseddeepvisualodometry] settings. DeepVO [wang2017deepvo] uses recurrent networks for temporal modeling, while SfMLearner [alisha2019unsupervised_learning_of_depth_and_ego_motion] jointly learns depth and motion without labels. DPVO [teed2023dpvo], inspired by RAFT [zachary2020raft], combines flow, confidence, and geometry with iterative GRU updates, while LeapVO [chen2024leapvo] increases robustness in dynamic scenes via keypoints. While being great at optimizing the rotations via bundle adjustment, these methods fail to estimate the scale of the scene, making their applicability in real-world scenarios limited. Moreover, they require camera parameters and are slow, which makes their deployment difficult. In contrast, VoT directly regresses relative camera poses in metric space and does not require camera parameters.
@@
Prior work introduced TSFormer [francani2025tsformer], which predicts camera poses directly from video sequences using an end-to-end approach. However, its performance remains significantly below that of hand-crafted methods [teed2023dpvo, chen2024leapvo] due to architecture, small training dataset, rotation representation, and inference scheme. Existing end-to-end methods struggle to perform in real-world settings because they rely on small, hard-to-scale architectures, and the gap between them and traditional methods remains substantial. By comparison, VoT uses a scalable end-to-end design which provided enough data to outperform traditional methods. When the data is limited, our method significantly reduces the performance gap between end-to-end and optimization-based pipelines, while being significantly faster.
@@
Video transformers. Since their introduction in machine translation [vaswani2017transformer], transformers have become the dominant architecture in NLP [devlin2019bert, openai2023gpt4] and vision [he2022mae, oquab2024dinov2]. Central to this framework, self-attention aggregates information across entire sequences, but its quadratic complexity makes direct application to videos inefficient. To mitigate this, prior work has proposed factorizing 3D attention into spatial and temporal components [zhang2021vidtr] or employing efficient variants such as divided space-time, sparse local-global, and axial attention [bertasius2021timesformer, arnab2021vivit, zhao2022tuber]. In this work, we adopt the principle of efficient video learning by decoupling self-attention into separate temporal and spatial attention components.
Large 3D models. A trend in 3D vision is the development of large-scale models that jointly estimate camera poses and dense geometry [wang20243d_reconstruction_with_spatial_memory, wang2025vggt, wang2025cut3r], or learn versatile representations for downstream tasks [wang2024dust3r, leroy2024mast3r]. These models are increasingly viewed as foundational, yet despite strong task-agnostic results they suffer from drift on long video sequences and produce scale-ambiguous poses. In contrast, VoT focuses solely on the task of camera pose estimation, showing improved accuracy and less drift. Moreover, our model is significantly faster due to its size.
@@ 3 Method
In this section, we describe the architecture of VoT; see Fig. 2 for its overview. It consists of three main components: a pre-trained encoder to extract a compact representation of each frame, a transformer-based decoder, and a small feed-forward network that makes the final pose predictions. Unlike previous methods such as DVPO [teed2023dpvo] or LeapVO [chen2024leapvo], VoT can be implemented in any deep learning framework, thanks to its streamlined architecture. Following the spirit of end-to-end frameworks [carion2020detr], our inference code can be written purely in PyTorch, without additional complexities. We believe that the simplicity and accessibility of our approach will encourage further participation from the research community.
@@ 3.1 Architecture
Pre-trained encoder. We utilize pre-trained networks commonly employed in previous studies [wang2024dust3r, wang2025cut3r] for encoding and extracting features from video frames. As the Vision Transformer (ViT) [dosovitskiy2020vit] has become the dominant architecture in computer vision, our pre-trained encoders are based on ViT. Specifically, we denote the sequence of input video frames as $V\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ is the number of frames in the video and $H,W$ denote the resolution of each frame. Each input image is then tokenized into $(h\cdot w)$ non-overlapping patches with $h{=}\frac{H}{p},w{=}\frac{W}{p}$ and $p$ as patch size. These tokens are processed through a series of ViT layers, producing image features $F\in\mathbb{R}^{T\times(h\cdot w)\times d}$, where $d$ is the hidden dimension of the ViT. The ViT layer follows the standard architecture, consisting of a multi-head self-attention module and a feed-forward network. Since the transformer architecture is permutation-invariant, we augment it with sinusoidal position encodings [vaswani2017transformer].
@@
$\begin{split}\hat{Q}=\hat{K}=\hat{V}&=F\text{[:, 1:, :]}^{\top},\\ \hat{Q},\hat{K},\hat{V}&\in\mathbb{R}^{(h\cdot w)\times T\times d}\end{split}$ (1)
$\begin{split}\hat{\mathrm{head}}_{i}&=\mathrm{Attention}(\hat{Q}W^{\hat{Q}}_{i},\hat{K}W^{\hat{K}}_{i},\hat{V}W^{\hat{V}}_{i}),\\ \hat{\mathrm{head}}_{i}&\in\mathbb{R}^{(h\cdot w)\times T\times d_{h}},\end{split}$ (2)
@@
$\begin{split}\hat{F}&=\mathrm{Concat}(\hat{\mathrm{head}}_{i},\ldots,\hat{\mathrm{head}}_{h})^{\top}W^{\hat{O}},\\ \hat{F}&\in\mathbb{R}^{T\times(h\cdot w)\times d}\end{split}$ (3)
$\begin{split}\mathrm{TemporalAttention}(\hat{Q},\hat{K},\hat{V})&=[F\text{[:,:1,:]},\hat{F}],\\ \mathrm{TemporalAttention}&\in\mathbb{R}^{T\times(h\cdot w+1)\times d}\end{split}$ (4)
@@
$\bar{Q}=\bar{K}=\bar{V}=\hat{F}\in\mathbb{R}^{T\times(h\cdot w+1)\times d_{h}}\\$ (5)
$\begin{split}\bar{\mathrm{head}}_{i}&=\mathrm{Attention}(\bar{Q}W^{\bar{Q}}_{i},\bar{K}W^{\bar{K}}_{i},\bar{V}W^{\bar{V}}_{i})\\ &\in\mathbb{R}^{T\times(h\cdot w+1)\times d_{h}},\end{split}$ (6)
@@
$\begin{split}\mathrm{SpatialAttention}(\bar{Q},\bar{K},\bar{V})&=\mathrm{Concat}(\bar{\mathrm{head}}_{i},\ldots,\bar{\mathrm{head}}_{h})W^{\bar{O}}\\ &\in\mathbb{R}^{T\times(h\cdot w+1)\times d}\end{split}$ (7)
@@
Ultimately, the camera embedding, $e\in\mathbb{R}^{T\times d}$, will be utilized to predict the camera poses for each frame.
@@ 3.2 Camera Poses Prediction
To regress the relative camera poses, the regression head takes a camera embedding from each frame as its input. A single linear layer then projects the embeddings to the rotation $\mathbb{F}_{R}\in\mathbb{R}^{(T-1)\times 9}$ and translation $\mathbb{F}_{t}\in\mathbb{R}^{(T-1)\times 3}$ vectors. Rotation vectors are projected to the closest valid rotation matrix by solving the special orthogonal Procrustes [bregier2021deepregression] problem:
@@
The optimization problem is solved via singular value decomposition of $\mathbb{F}_{R}$ following [umeyama1991lse]. $\mathbb{F}_{t}\in\mathbb{R}^{(T-1)\times 3}$ is used as a translation prediction directly.